# SyncDiff: Synchronized Motion Diffusion for Multi-Body Human-Object Interaction Synthesis

### v2

@@ A.2 Complete Proof of Our Two Synchronization Mechanisms
where $\lambda_{k}$ is a hyper-parameter, $\{(a_{k},b_{k},c_{k})\}_{k=1}^{|S|}$ is the set consisting of all triplets from $\hat{x}^{{}^{\prime}}$, where $c_{k}$ can be computed by $a_{k}$ and $b_{k}$ through combination operation or relative operation. For example, if $a_{k}=\hat{x}^{{}^{\prime}}_{u,o_{1}}$ and $b_{k}=\hat{x}^{{}^{\prime}}_{u,o_{2}}$, then $c_{k}=\hat{x}^{{}^{\prime}}_{b,o_{2}\rightarrow o_{1}}$, and $a_{k}\circ b_{k}=\text{rel}\left(\hat{x}^{{}^{\prime}}_{u,o_{1}},\hat{x}^{{}^{\prime}}_{u,o_{2}}\right)$. If $a_{k}=\hat{x}^{{}^{\prime}}_{u,o_{1}}$ and $b_{k}=\hat{x}^{{}^{\prime}}_{b,h_{1}\rightarrow o_{1}}$, then $c_{k}=\hat{x}^{{}^{\prime}}_{u,h_{1}}$, and $a_{k}\circ b_{k}=\text{comb}\left(\hat{x}^{{}^{\prime}}_{u,o_{1}},\hat{x}^{{}^{\prime}}_{b,h_{1}\rightarrow o_{1}}\right)$. The alignment likelihoods encompass the likelihoods of all binary computational relationships. Note that the alignment negative log-likelihood here is different from alignment loss $\mathcal{L}_{\text{align}}$ in the main text, which only considers relative operations. The intrinsic mathematical meaning of one item $\lambda_{k}\|c_{k}-a_{k}\circ b_{k}\|^{2}$ is similar to the above derivations, where we let $c_{k}$ follow a Gaussian distribution with a mean of $\hat{\mu}=a_{k}\circ b_{k}$ and a variance of $\sigma^{2}=\frac{1}{2\lambda_{k}}$.
@@
Now consider r fixing $c_{k}=\hat{x}^{{}^{\prime\prime}}$. It might be computed by $\kappa$ pairs of $(a_{k},b_{k})$. Take $\hat{x}^{{}^{\prime\prime}}=\hat{x}^{{}^{\prime}}_{u,o_{1}}$ as an example. Here $\kappa=m-1$, and $(a_{1},b_{1})=(\hat{x}^{{}^{\prime}}_{u,o_{2}},\hat{x}^{{}^{\prime}}_{b,o_{2}\rightarrow o_{1}})$, $(a_{2},b_{2})=(\hat{x}^{{}^{\prime}}_{u,o_{3}},\hat{x}^{{}^{\prime}}_{b,o_{3}\rightarrow o_{1}})$, $\dots$, $(a_{m-1},b_{m-1})=(\hat{x}^{{}^{\prime}}_{u,o_{m}},\hat{x}^{{}^{\prime}}_{b,o_{m}\rightarrow o_{1}})$. We need to simultaneously make $\hat{x}^{{}^{\prime\prime}}$ as close as possible to the corresponding part $\hat{\mu}^{{}^{\prime\prime}}=\hat{\mu}_{u,o_{1}}$ in $\hat{\mu}$ predicted by the diffusion model, while ensuring that $\hat{x}^{{}^{\prime\prime}}$ aligns with each predicted pair $(a_{k},b_{k})$. Add these terms together, maximizing total likelihood (the combination of data sample likelihood and alignment likelihood) is equivalent to minimizing
$\mathcal{P}_{\hat{x}^{{}^{\prime\prime}}}=\frac{1}{2\sigma^{2}}\|\hat{x}^{{}^{\prime\prime}}-\hat{\mu}^{{}^{\prime\prime}}\|_{2}^{2}+\sum\limits_{k=1}^{\kappa}\lambda_{k}\|\hat{x}^{{}^{\prime\prime}}-a_{k}\circ b_{k}\|_{2}^{2}$
@@
A problem here is that $\hat{\mu}$ is predicted by the model based on the result of step $t$, but $(a_{1},b_{1})$, $\dots$, $(a_{\kappa},b_{\kappa})$ all belong to step $(t-1)$ along with $\hat{x}^{{}^{\prime\prime}}$. Here, we make an assumption that $\hat{x}_{t}$ and $\hat{x}_{t-1}$ are close, so that we can take $(a_{1},b_{1})$, $(a_{2},b_{2})$, $\dots$, $(a_{\kappa},b_{\kappa})$ from $\hat{x}_{t}=\hat{x}$.
@@
Denote $\hat{\mu}^{{}^{\prime\prime}}$ as $f_{0}$, and $a_{1}\circ b_{1}$, $a_{2}\circ b_{2}$, $\dots$, $a_{\kappa}\circ b_{\kappa}$ as $f_{1}$, $f_{2}$, $\dots$, $f_{\kappa}$. Here $f_{0}$, $f_{1}$, $\dots$, $f_{\kappa}$ are all deterministic values calculated from some certain parts of $\hat{x}_{t}$. Also let $\lambda_{0}=\frac{1}{2\sigma^{2}}$, then
$\mathcal{P}_{\hat{x}^{\prime\prime}}$ $=\sum\limits_{k=0}^{\kappa}\lambda_{k}\|\hat{x}^{\prime\prime}-f_{k}\|_{2}^{2}$
$=\sum\limits_{k=0}^{\kappa}\lambda_{k}\left(\hat{x}^{\prime\prime\top}\hat{x}^{\prime\prime}-2f_{k}^{\top}\hat{x}^{\prime\prime}+f_{k}^{\top}f_{k}\right)$
$=\left(\sum\limits_{k=0}^{\kappa}\lambda_{k}\right)\|\hat{x}^{\prime\prime}\|_{2}^{2}-2\left(\sum\limits_{k=0}^{\kappa}\lambda_{k}f_{k}\right)^{\top}\hat{x}^{\prime\prime}+\sum\limits_{k=0}^{\kappa}\lambda_{k}\|f_{k}\|_{2}^{2}$
@@
$\hat{\mu}^{\prime}=\sum\limits_{k=0}^{\kappa}\frac{\lambda_{k}}{\sum\limits_{k=0}^{\kappa}\lambda_{k}}f_{k}$
$\sigma^{\prime 2}=\frac{1}{2\left(\sum\limits_{k=0}^{\kappa}\lambda_{k}\right)}$
@@
1. For individual motions of rigid body $o_{j}(j\in[1,m])$ (Here we assume that $m>1$, otherwise there is no need for explicit synchronization on this part), relevant pairs of $(a_{k},b_{k})$ consist of $(\hat{x}_{u,o_{j^{{}^{\prime}}}},\hat{x}_{b,o_{j}\rightarrow o_{j^{{}^{\prime}}}})(j^{{}^{\prime}}\neq j)$. $\lambda_{0}=\frac{1}{2\sigma^{2}}$, $\lambda_{1}=\lambda_{2}=\dots=\lambda_{m-1}=\frac{\overline{\lambda}}{m-1}$. Here $\overline{\lambda}$ is an empirical value, satisfying
$\overline{\lambda}=\frac{\lambda_{\text{exp}}}{\mathcal{R}}\sum\limits_{r=1}^{\mathcal{R}}\frac{1}{2\sigma_{i_{r}}^{2}}$
where $1\leq i_{1}<i_{2}<\dots<i_{\mathcal{R}}\leq\mathcal{T}$ are the synchronization timesteps, and $\sigma_{i_{1}}$, $\sigma_{i_{2}}$, $\dots$, $\sigma_{i_{\mathcal{R}}}$ are the original correspondent standard variances(without synchronization). The value of hyperparameter $\lambda_{\text{exp}}$ can be found in Table  10. Therefore,
$\hat{\mu}^{{}^{\prime}}_{u,o_{j}}$
$=\frac{1}{\frac{1}{2\sigma^{2}}+\overline{\lambda}}\left(\lambda_{0}\hat{\mu}_{u,o_{j}}+\sum\limits_{j^{{}^{\prime}}\neq j}\frac{\overline{\lambda}}{m-1}\text{comb}\left(\hat{x}_{u,o_{j^{{}^{\prime}}}},\hat{x}_{b,o_{j}\rightarrow o_{j^{{}^{\prime}}}}\right)\right)$
$=\frac{1}{1+2\sigma^{2}\overline{\lambda}}\hat{\mu}_{u,o_{j}}+\frac{\frac{2}{m-1}\sigma^{2}\overline{\lambda}}{1+2\sigma^{2}\overline{\lambda}}\sum\limits_{j^{{}^{\prime}}\neq j}\text{comb}\left(\hat{x}_{u,o_{j^{{}^{\prime}}}},\hat{x}_{b,o_{j}\rightarrow o_{j^{{}^{\prime}}}}\right)$
2. For individual motions of articulated skeleton $h_{i}(i\in[1,n])$, relevant pairs of $(a_{k},b_{k})$ consist of $(\hat{x}_{u,o_{j}},\hat{x}_{b,h_{i}\rightarrow o_{j}})(j\in[1,m])$. $\lambda_{0}=\frac{1}{2\sigma^{2}}$, $\lambda_{1}=\lambda_{2}=\dots=\lambda_{m}=\frac{\overline{\lambda}}{m}$. Therefore,
$\hat{\mu}^{{}^{\prime}}_{u,h_{i}}$ $=\frac{1}{\frac{1}{2\sigma^{2}}+\overline{\lambda}}\left(\lambda_{0}\hat{\mu}_{u,h_{i}}+\sum\limits_{j\in[1,m]}\frac{\overline{\lambda}}{m}\text{comb}\left(\hat{x}_{u,o_{j}},\hat{x}_{b,h_{i}\rightarrow o_{j}}\right)\right)$
$=\frac{1}{1+2\sigma^{2}\overline{\lambda}}\hat{\mu}_{u,o_{j}}+\frac{\frac{2}{m}\sigma^{2}\overline{\lambda}}{1+2\sigma^{2}\overline{\lambda}}\sum\limits_{j\in[1,m]}\text{comb}\left(\hat{x}_{u,o_{j}},\hat{x}_{b,h_{i}\rightarrow o_{j}}\right)$
@@
$\hat{\mu}^{{}^{\prime}}_{b,o_{j}\rightarrow o_{j^{{}^{\prime}}}}$ $=\frac{1}{\frac{1}{2\sigma^{2}}+\overline{\lambda}}\left(\lambda_{0}\hat{\mu}_{b,o_{j}\rightarrow o_{j^{{}^{\prime}}}}+\lambda_{1}\text{rel}\left(\hat{x}_{u,o_{j^{{}^{\prime}}}},\hat{x}_{u,o_{j}}\right)\right)$
$=\frac{1}{1+2\sigma^{2}\overline{\lambda}}\hat{\mu}_{b,o_{j}\rightarrow o_{j^{{}^{\prime}}}}+\frac{2\sigma^{2}\overline{\lambda}}{1+2\sigma^{2}\overline{\lambda}}\text{rel}\left(\hat{x}_{u,o_{j^{{}^{\prime}}}},\hat{x}_{u,o_{j}}\right)$
$\hat{\mu}^{{}^{\prime}}_{b,h_{i}\rightarrow o_{j}}$ $=\frac{1}{\frac{1}{2\sigma^{2}}+\overline{\lambda}}\left(\lambda_{0}\hat{\mu}_{b,h_{i}\rightarrow o_{j}}+\lambda_{1}\text{rel}\left(\hat{x}_{u,o_{j}},\hat{x}_{u,h_{i}}\right)\right)$
$=\frac{1}{1+2\sigma^{2}\overline{\lambda}}\hat{\mu}_{b,h_{i}\rightarrow o_{j}}+\frac{2\sigma^{2}\overline{\lambda}}{1+2\sigma^{2}\overline{\lambda}}\text{rel}\left(\hat{x}_{u,o_{j}},\hat{x}_{u,h_{i}}\right)$
@@ A.3 Algorithm for Explicit Synchronization
2:     $\lambda\leftarrow 0$, $\mathcal{R}\leftarrow 0$
3:     for $t\leftarrow\mathcal{T}$ to $1$ do
4:         if $t\bmod t_{0}=\lfloor t_{0}/2\rfloor$ then
5:              $\lambda\leftarrow\lambda+\frac{1}{2\sigma_{t}^{2}}$, $\mathcal{R}\leftarrow\mathcal{R}+1$
@@
8:     $\overline{\lambda}\leftarrow\lambda_{\text{exp}}\cdot\frac{\lambda}{\mathcal{R}}$
9:     $\hat{x}_{\mathcal{T}}\sim\mathcal{N}(0,I)$
10:     for $t\leftarrow\mathcal{T}$ to $1$ do
@@
12:         if $t\bmod t_{0}=\lfloor t_{0}/2\rfloor$ then
@@
17:                       $\hat{x}^{{}^{\prime}}_{u,o_{j}}\leftarrow\hat{x}^{{}^{\prime}}_{u,o_{j}}+\lambda_{1}\cdot\hat{\mu}_{u,o_{j}}$
@@
20:                           if $j>j^{{}^{\prime}}$ then
21:                                $\hat{x}^{{}^{\prime}}_{u,o_{j}}\leftarrow\hat{x}^{{}^{\prime}}_{u,o_{j}}+\frac{\lambda_{1}}{m-1}\cdot\text{comb}\left(\hat{x}_{u,o_{j^{{}^{\prime}}}},\hat{x}_{b,o_{j}\rightarrow o_{j^{{}^{\prime}}}}\right)$
22:                                $\hat{x}^{{}^{\prime}}_{b,o_{j}\rightarrow o_{j^{{}^{\prime}}}}\leftarrow\hat{x}^{{}^{\prime}}_{b,o_{j}\rightarrow o_{j^{{}^{\prime}}}}+\lambda_{1}\cdot\text{rel}\left(\hat{x}_{u,o_{j^{{}^{\prime}}}},\hat{x}_{u,o_{j}}\right)$
23:                           else
24:                                $\hat{x}^{{}^{\prime}}_{u,o_{j}}\leftarrow\hat{x}^{{}^{\prime}}_{u,o_{j}}+\frac{\lambda_{1}}{m-1}\cdot\text{comb}\left(\hat{x}_{u,o_{j^{{}^{\prime}}}},\hat{x}_{b,o_{j^{{}^{\prime}}}\rightarrow o_{j}}\right)$
25:                           end if
26:                       end for
27:                  end if
28:              end for
29:              for $i\leftarrow 1$ to $n$ do
30:                  for $j\leftarrow 1$ to $m$ do
31:                       $\hat{x}^{{}^{\prime}}_{u,h_{i}}\leftarrow\hat{x}^{{}^{\prime}}_{u,h_{i}}+\frac{\lambda_{1}}{m}\cdot\text{comb}\left(\hat{x}_{u,o_{j}},\hat{x}_{b,h_{i}\rightarrow o_{j}}\right)$
32:                       $\hat{x}^{{}^{\prime}}_{b,h_{i}\rightarrow o_{j}}\leftarrow\hat{x}^{{}^{\prime}}_{b,h_{i}\rightarrow o_{j}}+\lambda_{1}\cdot\text{rel}\left(\hat{x}_{u,o_{j}},\hat{x}_{u,h_{i}}\right)$
33:                  end for
34:              end for
35:              $\sigma^{{}^{\prime}}\leftarrow\sqrt{\frac{\sigma_{t}^{2}}{1+2\sigma_{t}^{2}\overline{\lambda}}}$
36:         else
37:              $\hat{x}^{{}^{\prime}}\leftarrow\hat{\mu}$
38:              $\sigma^{{}^{\prime}}\leftarrow\sigma_{t}$
39:         end if
40:         $\epsilon\sim\mathcal{N}(0,I)$
41:         $\hat{x}^{{}^{\prime}}\leftarrow\hat{x}^{{}^{\prime}}+\sigma^{{}^{\prime}}\cdot\epsilon$
42:         $\hat{x}_{t-1}\leftarrow\hat{x}^{{}^{\prime}}$
43:     end for
44:end procedure
Appendix B More Details about Data Processing and Model Architecture
B.1 Hyperparameters for Frequency-based Decomposition and Explicit Synchronization
In motion sequence decomposition, we discard signals with frequencies higher than $L$. We analyzed the impact of different $L$ on the maximum error $\epsilon$, which is averaged on the dimension of $t$ and takes the maximum across all degrees of freedom. The errors listed in Table  5 are averaged over all data samples in the datasets and are measured in millimeters. In practice, we take $L=16$.
$L$ TACO CORE4D OAKINK2 GRAB
6 8.3 33.1 6.1 8.4
8 6.8 15.3 4.9 6.9
12 4.9 16.4 3.7 5.1
16 3.9 11.4 3.0 4.0
20 3.2 9.0 2.6 3.2
25 2.6 7.2 2.2 2.6

### v3

@@ A.2 Complete Proof of Our Two Synchronization Mechanisms
where $\lambda_{v\in[1,V]}$ are a hyperparameters, $\{(a_{v},b_{v},c_{v})\}_{v=1}^{|V|}$ is the set consisting of all triplets $(\hat{x}^{{}^{\prime}}_{\textbf{b}_{1}},\hat{x}^{{}^{\prime}}_{\textbf{b}_{2}},\hat{x}^{{}^{\prime}}_{\textbf{b}_{2}\rightarrow\textbf{b}_{1}})$, where $c_{v}$ can be computed by $a_{v}$ and $b_{v}$ through combination operation or relative operation. For example, if $a_{v}=\hat{x}^{{}^{\prime}}_{o_{1}}$ and $b_{v}=\hat{x}^{{}^{\prime}}_{o_{2}}$, then $c_{v}=\hat{x}^{{}^{\prime}}_{o_{2}\rightarrow o_{1}}$, and $a_{v}\circ b_{v}=\text{rel}\left(\hat{x}^{{}^{\prime}}_{o_{1}},\hat{x}^{{}^{\prime}}_{o_{2}}\right)$. If $a_{v}=\hat{x}^{{}^{\prime}}_{o_{1}}$ and $b_{v}=\hat{x}^{{}^{\prime}}_{h_{1}\rightarrow o_{1}}$, then $c_{v}=\hat{x}^{{}^{\prime}}_{h_{1}}$, and $a_{v}\circ b_{v}=\text{comb}\left(\hat{x}^{{}^{\prime}}_{o_{1}},\hat{x}^{{}^{\prime}}_{h_{1}\rightarrow o_{1}}\right)$. The alignment likelihoods encompass the likelihoods of all binary computational relationships. Note that the alignment negative log-likelihood here is different from alignment loss $\mathcal{L}_{\text{align}}$ in the main text, which only considers relative operations rel, without combination operations comb. The intrinsic mathematical meaning of one item $\lambda_{v}\|c_{v}-a_{v}\circ b_{v}\|^{2}$ is similar to the above derivations of $\mathcal{L}_{\text{align}}$, where we let $c_{v}$ follow a Gaussian distribution with a mean of $\hat{\mu}=a_{v}\circ b_{v}$ and a variance of $\sigma^{2}=\frac{1}{2\lambda_{v}}$.
@@
Now consider fixing some $c_{v}=\hat{x}^{{}^{\prime\prime}}$ (This is the motion of one single body, which is different from $\hat{x}$ and $\hat{x}^{{}^{\prime}}$, the representation comprising of all individual/relative motions). It might be computed by $K$ pairs of $(a_{k},b_{k})$. Take $\hat{x}^{{}^{\prime\prime}}=\hat{x}^{{}^{\prime}}_{o_{1}}$ as an example. Here $K=m-1$, and $(a_{1},b_{1})=(\hat{x}^{{}^{\prime}}_{o_{2}},\hat{x}^{{}^{\prime}}_{o_{1}\rightarrow o_{2}})$, $(a_{2},b_{2})=(\hat{x}^{{}^{\prime}}_{o_{3}},\hat{x}^{{}^{\prime}}_{o_{1}\rightarrow o_{3}})$, $\dots$, $(a_{m-1},b_{m-1})=(\hat{x}^{{}^{\prime}}_{o_{m}},\hat{x}^{{}^{\prime}}_{o_{1}\rightarrow o_{m}})$. We need to simultaneously make $\hat{x}^{{}^{\prime\prime}}$ as close as possible to the corresponding part $\hat{\mu}^{{}^{\prime\prime}}=\hat{\mu}_{o_{1}}$ in $\hat{\mu}$ predicted by the diffusion model, while ensuring that $\hat{x}^{{}^{\prime\prime}}$ aligns with each predicted pair $(a_{k},b_{k})$. Add these terms together, maximizing total likelihood (the combination of data sample likelihood and alignment likelihoods) is equivalent to minimizing
$\mathcal{P}_{\hat{x}^{{}^{\prime\prime}}}=\frac{1}{2\sigma^{2}}\|\hat{x}^{{}^{\prime\prime}}-\hat{\mu}^{{}^{\prime\prime}}\|_{2}^{2}+\sum\limits_{k=1}^{K}\lambda_{k}\|\hat{x}^{{}^{\prime\prime}}-a_{k}\circ b_{k}\|_{2}^{2}$
@@
A problem here is that $\hat{\mu}$ is predicted by the model based on the result of step $t$, but $(a_{1},b_{1})$, $\dots$, $(a_{K},b_{K})$ all belong to step $(t-1)$ along with $\hat{x}^{{}^{\prime\prime}}$. Here, we make an assumption that $\hat{x}_{t}$ and $\hat{x}_{t-1}$ are close, so that we can take $(a_{1},b_{1})$, $(a_{2},b_{2})$, $\dots$, $(a_{K},b_{K})$ from $\hat{x}_{t}=\hat{x}$.
@@
Denote $\hat{\mu}^{{}^{\prime\prime}}$ as $f_{0}$, and $a_{1}\circ b_{1}$, $a_{2}\circ b_{2}$, $\dots$, $a_{K}\circ b_{K}$ as $f_{1}$, $f_{2}$, $\dots$, $f_{K}$. Here $f_{0}$, $f_{1}$, $\dots$, $f_{K}$ are all deterministic values calculated from some certain parts of $\hat{x}_{t}$. Also let $\lambda_{0}=\frac{1}{2\sigma^{2}}$, then
$\mathcal{P}_{\hat{x}^{\prime\prime}}$ $=\sum\limits_{k=0}^{K}\lambda_{k}\|\hat{x}^{\prime\prime}-f_{k}\|_{2}^{2}$
$=\sum\limits_{k=0}^{K}\lambda_{k}\left(\hat{x}^{\prime\prime\top}\hat{x}^{\prime\prime}-2f_{k}^{\top}\hat{x}^{\prime\prime}+f_{k}^{\top}f_{k}\right)$
$=\left(\sum\limits_{k=0}^{K}\lambda_{k}\right)\|\hat{x}^{\prime\prime}\|_{2}^{2}-2\left(\sum\limits_{k=0}^{K}\lambda_{k}f_{k}\right)^{\top}\hat{x}^{\prime\prime}+\sum\limits_{k=0}^{K}\lambda_{k}\|f_{k}\|_{2}^{2}$
@@
$\hat{\mu}^{\prime}=\sum\limits_{k=0}^{K}\frac{\lambda_{k}}{\sum\limits_{k=0}^{K}\lambda_{k}}f_{k}$
$\sigma^{\prime 2}=\frac{1}{2\left(\sum\limits_{k=0}^{K}\lambda_{k}\right)}$
@@
1. For individual motions of rigid body $o_{j}(j\in[1,m])$ (Here we assume that $m>1$, otherwise there is no need for explicit synchronization on this part), relevant pairs of $(a_{k},b_{k})$ consist of $(\hat{x}_{o_{j^{{}^{\prime}}}},\hat{x}_{o_{j}\rightarrow o_{j^{{}^{\prime}}}})(j^{{}^{\prime}}\neq j)$. $\lambda_{0}=\frac{1}{2\sigma^{2}}$, $\lambda_{1}=\lambda_{2}=\dots=\lambda_{m-1}=\frac{\overline{\lambda}}{m-1}$. Here $\overline{\lambda}$ is an empirical value, satisfying
$\overline{\lambda}=\frac{\lambda_{\text{exp}}}{R}\sum\limits_{r=1}^{R}\frac{1}{2\sigma_{t_{r}}^{2}}$
where $1\leq t_{1}<t_{2}<\dots<t_{R}\leq T$ are the synchronization timesteps, and $\sigma_{t_{1}}$, $\sigma_{t_{2}}$, $\dots$, $\sigma_{t_{R}}$ are the original correspondent standard variances(without synchronization). The value of hyperparameter $\lambda_{\text{exp}}$ can be found in Table  17. Therefore,
$\hat{\mu}^{{}^{\prime}}_{o_{j}}$
$=\frac{1}{\frac{1}{2\sigma^{2}}+\overline{\lambda}}\left(\lambda_{0}\hat{\mu}_{o_{j}}+\sum\limits_{j^{{}^{\prime}}\neq j}\frac{\overline{\lambda}}{m-1}\text{comb}\left(\hat{x}_{o_{j^{{}^{\prime}}}},\hat{x}_{o_{j}\rightarrow o_{j^{{}^{\prime}}}}\right)\right)$
$=\frac{1}{1+2\sigma^{2}\overline{\lambda}}\hat{\mu}_{o_{j}}+\frac{\frac{2}{m-1}\sigma^{2}\overline{\lambda}}{1+2\sigma^{2}\overline{\lambda}}\sum\limits_{j^{{}^{\prime}}\neq j}\text{comb}\left(\hat{x}_{o_{j^{{}^{\prime}}}},\hat{x}_{o_{j}\rightarrow o_{j^{{}^{\prime}}}}\right)$
2. For individual motions of articulated skeleton $h_{i}(i\in[1,n])$, relevant pairs of $(a_{k},b_{k})$ consist of $(\hat{x}_{o_{j}},\hat{x}_{h_{i}\rightarrow o_{j}})(j\in[1,m])$. $\lambda_{0}=\frac{1}{2\sigma^{2}}$, $\lambda_{1}=\lambda_{2}=\dots=\lambda_{m}=\frac{\overline{\lambda}}{m}$. Therefore,
$\hat{\mu}^{{}^{\prime}}_{h_{i}}$ $=\frac{1}{\frac{1}{2\sigma^{2}}+\overline{\lambda}}\left(\lambda_{0}\hat{\mu}_{h_{i}}+\sum\limits_{j\in[1,m]}\frac{\overline{\lambda}}{m}\text{comb}\left(\hat{x}_{o_{j}},\hat{x}_{h_{i}\rightarrow o_{j}}\right)\right)$
$=\frac{1}{1+2\sigma^{2}\overline{\lambda}}\hat{\mu}_{o_{j}}+\frac{\frac{2}{m}\sigma^{2}\overline{\lambda}}{1+2\sigma^{2}\overline{\lambda}}\sum\limits_{j\in[1,m]}\text{comb}\left(\hat{x}_{o_{j}},\hat{x}_{h_{i}\rightarrow o_{j}}\right)$
@@
$\hat{\mu}^{{}^{\prime}}_{o_{j}\rightarrow o_{j^{{}^{\prime}}}}$ $=\frac{1}{\frac{1}{2\sigma^{2}}+\overline{\lambda}}\left(\lambda_{0}\hat{\mu}_{o_{j}\rightarrow o_{j^{{}^{\prime}}}}+\lambda_{1}\text{rel}\left(\hat{x}_{o_{j^{{}^{\prime}}}},\hat{x}_{o_{j}}\right)\right)$
$=\frac{1}{1+2\sigma^{2}\overline{\lambda}}\hat{\mu}_{o_{j}\rightarrow o_{j^{{}^{\prime}}}}+\frac{2\sigma^{2}\overline{\lambda}}{1+2\sigma^{2}\overline{\lambda}}\text{rel}\left(\hat{x}_{o_{j^{{}^{\prime}}}},\hat{x}_{o_{j}}\right)$
$\hat{\mu}^{{}^{\prime}}_{h_{i}\rightarrow o_{j}}$ $=\frac{1}{\frac{1}{2\sigma^{2}}+\overline{\lambda}}\left(\lambda_{0}\hat{\mu}_{h_{i}\rightarrow o_{j}}+\lambda_{1}\text{rel}\left(\hat{x}_{o_{j}},\hat{x}_{h_{i}}\right)\right)$
$=\frac{1}{1+2\sigma^{2}\overline{\lambda}}\hat{\mu}_{h_{i}\rightarrow o_{j}}+\frac{2\sigma^{2}\overline{\lambda}}{1+2\sigma^{2}\overline{\lambda}}\text{rel}\left(\hat{x}_{o_{j}},\hat{x}_{h_{i}}\right)$
@@ A.3 Algorithm for Explicit Synchronization
2:     $\lambda\leftarrow 0$, $R\leftarrow 0$
3:     for $t\leftarrow T$ to $1$ do
4:         if $t\bmod s=\lfloor s/2\rfloor$ then
5:              $\lambda\leftarrow\lambda+\frac{1}{2\sigma_{t}^{2}}$, $R\leftarrow R+1$
@@
8:     $\overline{\lambda}\leftarrow\lambda_{\text{exp}}\cdot\frac{\lambda}{R}$
9:     $\hat{x}_{T}\sim\mathcal{N}(0,I)$
10:     for $t\leftarrow T$ to $1$ do
@@
12:         if $t\bmod s=\lfloor s/2\rfloor$ then
@@
17:                       $\hat{x}^{{}^{\prime}}_{o_{j}}\leftarrow\hat{x}^{{}^{\prime}}_{o_{j}}+\lambda_{1}\cdot\hat{\mu}_{o_{j}}$
@@
20:                           $\hat{x}^{{}^{\prime}}_{o_{j}}\leftarrow\hat{x}^{{}^{\prime}}_{o_{j}}+\frac{\lambda_{1}}{m-1}\cdot\text{comb}\left(\hat{x}_{o_{j^{{}^{\prime}}}},\hat{x}_{o_{j}\rightarrow o_{j^{{}^{\prime}}}}\right)$
21:                           $\hat{x}^{{}^{\prime}}_{o_{j}\rightarrow o_{j^{{}^{\prime}}}}\leftarrow\hat{x}^{{}^{\prime}}_{o_{j}\rightarrow o_{j^{{}^{\prime}}}}+\lambda_{1}\cdot\text{rel}\left(\hat{x}_{o_{j^{{}^{\prime}}}},\hat{x}_{o_{j}}\right)$
22:                       end for
23:                  end if
24:              end for
25:              for $i\leftarrow 1$ to $n$ do
26:                  for $j\leftarrow 1$ to $m$ do
27:                       $\hat{x}^{{}^{\prime}}_{h_{i}}\leftarrow\hat{x}^{{}^{\prime}}_{h_{i}}+\frac{\lambda_{1}}{m}\cdot\text{comb}\left(\hat{x}_{o_{j}},\hat{x}_{h_{i}\rightarrow o_{j}}\right)$
28:                       $\hat{x}^{{}^{\prime}}_{h_{i}\rightarrow o_{j}}\leftarrow\hat{x}^{{}^{\prime}}_{h_{i}\rightarrow o_{j}}+\lambda_{1}\cdot\text{rel}\left(\hat{x}_{o_{j}},\hat{x}_{h_{i}}\right)$
29:                  end for
30:              end for
31:              $\sigma^{{}^{\prime}}\leftarrow\sqrt{\frac{\sigma_{t}^{2}}{1+2\sigma_{t}^{2}\overline{\lambda}}}$
32:         else
33:              $\hat{x}^{{}^{\prime}}\leftarrow\hat{\mu}$
34:              $\sigma^{{}^{\prime}}\leftarrow\sigma_{t}$
35:         end if
36:         $\epsilon\sim\mathcal{N}(0,I)$
37:         $\hat{x}^{{}^{\prime}}\leftarrow\hat{x}^{{}^{\prime}}+\sigma^{{}^{\prime}}\cdot\epsilon$
38:         $\hat{x}_{t-1}\leftarrow\hat{x}^{{}^{\prime}}$
39:     end for
40:end procedure
Appendix B Supplementary Experiments
B.1 Determine the cutoff boundary $L$ for Frequency Decomposition
In frequency decomposition (Section  3.3), we discard signals with frequencies higher than $\phi_{L}/2\pi$, where $\phi_{L}=L/N$, and $L$ is the cutoff boundary. We analyzed the impact of different $L$ on the maximum error $\epsilon$, which is averaged on the dimension of $N$ (number of frames) and takes the maximum across all degrees of freedom (the dimension of $D_{\text{sum}}$). The errors listed in Table  6 are averaged over all data samples in the datasets and are measured in millimeters.
$L$ TACO CORE4D OAKINK2 GRAB CG-HOI
6 8.3 33.1 6.1 8.4 11.5
8 6.8 15.3 4.9 6.9 7.8
12 4.9 13.4 3.7 5.1 6.8
16 3.9 11.4 3.0 4.0 5.9
20 3.2 9.0 2.6 3.2 5.2
25 2.6 7.2 2.2 2.6 4.8